A pharmacist has been monitoring sales of a certain over-the-counter pain reliever. Daily sales during the last 15 days were Day: 1 2 3 4 5 6 7 8 9 Number sold: 33 35 40 42 42 42 43 48 49 Day: 10 11 12 13 14 15 Number sold: 46 52 46 48 57 50 a. Assume the data refer to demand rather than sales. Using trend-adjusted exponential smoothing with an initial forecast of 48 for Day 8, an initial trend estimate of 2, and α = β = .2, develop demand forecasts for Days 9 through 16. Then compute the resultant MSE using the error values from Days 8 through 15. (Round your intermediate period-by-period forecast and error values to 3 decimal places. Round your final MSE answer to 3 decimal places.)
Answer #1

Formulae to be used for trend-adjusted exponential smoothing:

TAFt+1 = St + Tt
St = TAFt + α (At - TAFt)
Tt = Tt-1 + β (TAFt - TAFt-1 - Tt-1)

Calculations